### v1

# Consistency in Language Models: Current Landscape, Challenges, and Future Directions

Abstract
The hallmark of effective language use lies in consistency—expressing similar meanings in similar contexts and avoiding contradictions. While human communication naturally demonstrates this principle, state-of-the-art language models struggle to maintain reliable consistency across different scenarios. This paper examines the landscape of consistency research in AI language systems, exploring both formal consistency (including logical rule adherence) and informal consistency (such as moral and factual coherence). We analyze current approaches to measure aspects of consistency, identify critical research gaps in standardization of definitions, multilingual assessment, and methods to improve consistency. Our findings point to an urgent need for robust benchmarks to measure and interdisciplinary approaches to ensure consistency in the application of language models on domain-specific tasks while preserving the utility and adaptability.
1 Introduction
Consistency—broadly defined as using language similarly in similar settings or avoiding contradictions when using language—is among the most important forms of generalization in the use of language. This ability to maintain consistent outputs is essential for building reliable AI systems that users can trust and depend on. Consistency serves as both a natural expectation users have when interacting with language technologies and a prerequisite for deploying these models in high-stakes domains (Elazar et al., 2021; Jang et al., 2022; Kim et al., 2025). However, the most advanced large language models (LLMs) struggle with this and demonstrate multiple examples of inconsistent behavior (Elazar et al., 2021; Raj et al., 2025). Although such examples have been documented in multiple studies, there are no standard approaches to assessing model consistency. As such, there is an ongoing risk of overestimating the performance of state-of-the-art models, as well as of underestimating the risks and potential harms elicited by such models.
There have been some early attempts to analyze the behavior of models in terms of consistency, to understand the roots of inconsistency, and to discover ways to mitigate inconsistency and/or to use inconsistent outputs to improve the model. However, research on language model consistency faces multiple challenges, starting from a lack of agreement on terminology and evaluation methods, followed by limitations related to data and models. In this paper, we present a review of current research on consistency in language models, highlight the most pressing challenges, and provide recommendations for future research directions. For greater specificity, we restrict our attention to text-only language models, on which an overwhelming majority of existing research on consistency has been based. See Appendix A for references to some research on multimodal consistency.
2 A Review of Consistency Research in ML
Consistency in language models connects to several critical areas in AI research: hallucinations (enerating made-up information that may contradict known facts, factuality (agreement with real-world knowledge), misinformation (false claims misleading users), and reasoning (logical coherence across statements). As such, consistency evaluation forms a crucial component of foundation model assessment frameworks. This review surveys literature on consistency in language models from 2019 to 2025, focusing on peer-reviewed publications from major AI conferences and influential preprints that explicitly address consistency through evaluation, enhancement techniques, or theoretical analysis.
2.1 Terminology and Consistency Types
The terminology used to describe the consistency of language models is often confusing, as there is not a single, commonly agreed-upon definition of consistency. Authors either come up with their own definition of the concept that aligns best with the specifics of the work they focus on, or use an overly broad definition, or sometimes just omit defining the term altogether. As a result, existing research studies present multiple narrowly focused definitions of consistency that often cover very different aspects of model behavior and sometimes even contradict each other.
Despite disagreements about consistency definitions in previous studies of language models, all of them, in one way or another, relate to the behavior of these models. Given this interest in the behavior of models and the implications for potential model applications, in this paper we limit the otherwise broader concept of consistency to behavioral consistency. Behavioral consistency has been thoroughly analyzed in the field of human personality psychology, where it refers to adhering to the same principles in behavioral patterns across a large variety of situations and sometimes over many years (Funder & Colvin, 1991). In psychology, behavioral consistency is tightly linked with the predictability of behavior, which is equally important for the applications of language models. Based on how behavioral consistency is approached in the literature, in this paper, we further categorize the different types of consistency into two large groups: logical/formal and non-logical/informal.
2.1.1 Logical / formal consistency
Logical consistency in relation to large language models was introduced by Jang et al. (2022), as an ability of the model to make decisions without logical contradiction. The rules and principles of formal logic are applied to assess the behavior of a model in a methodical way, making the evaluation standardized and easy to follow. Based on these principles, Jang et al. (2022) classify consistency into negational, symmetric, transitive, and additive types. Negational consistency follows the logical negation property ($p$ is true $\Leftrightarrow\neg{p}$ is false), i.e. LM’s predictions should be opposite for texts with the opposite meanings. Symmetric consistency follows the rule $f(x,y)=f(y,x)$ and implies that LM ’s predictions should be invariant to the input text swap. Transitive consistency can measure deductive reasoning ability and follows the property of transitive inference, represented as $X\rightarrow Y\land Y\rightarrow Z$ then $X\rightarrow Z$. This type of consistency was analyzed in NLI tasks Li et al. (2019) as well as in question answering Asai & Hajishirzi (2020b); Mitchell et al. (2022).
Semantic consistency, another subpart of the Jang et al. (2022) definition, is one of the most frequently used concepts in existing consistency research studies. The idea of semantic consistency is derived from the semantic equivalence property, represented as $f(X)=f(Y)$ if $X$ and $Y$ mean the same. It was explored by Elazar et al. (2021) as the ability of a model to make consistent decisions in semantically equivalent contexts, and a very similar understanding of the concept was later used in multiple studies Raj et al. (2023; 2022); Ohmer et al. (2024).
2.1.2 Nonlogical / informal consistency.
The nonlogical, or informal, type of consistency covers all the other definitions that do not follow the rules of formal logic. For example, Bonagiri et al. (2024) highlight the importance of moral consistency, as the ability to preserve non-contradictory moral values across different types of situations (Arvanitis and Kalliris, 2020; Marcus, 1980), as a part of LLM alignment. Their approach consists of generating semantically equivalent scenarios and employing consistency checks to see if a target LLM gets the same Semantic Graph Entropy (SaGE) score while responding to these scenarios. To assess the reliability of SaGE, they use human evaluation. Jain et al. 2024 investigate norm inconsistency in LLMs, which they define as condition where LLMs apply different norms in similar situations. They specifically focus on applications of models in high-risk domains.
Informational and/or factual consistency is another subpart of the Jang et al. (2022) definition and is a term frequently used in consistency research work. Manakul et al. (2023) use the term informational consistency, without explaining or defining it further, in developing a method of fact-checking the responses of black-box models. The term factual consistency is often used in the context of automatic summarization Wang et al. (2020). Factual inconsistency is often referred to as hallucinations and/or faithfulness, i.e. models generating new information that contradicts the source document (e.g. Tam et al, 2023, Maynez et al., 2020). Definitions of factual consistency are often not clearly specified and instead replaced with human annotations, as in e.g. Tam et al. (2023). [^1] There is a case to be made that factual and moral consistency are better construed as a logical property. For instance, factual consistency could be construed as the absence of contradictions between a model’s outputs, and factually true statements.
Some more recent work clarifies the concept of consistency. A recent study by Parcalabescu & Frank (2024) on natural language explanations contrasts faithfulness and self-consistency. Self-consistency examines whether similar inputs produce consistent explanations—essentially measuring explanation stability across input variations. Faithfulness, meanwhile, evaluates whether the explanation behind a certain model-generated answer accurately reflects the model’s reasoning process to come up with that answer. While related, they involve different evaluation approaches. Self-consistency requires testing multiple input variations (which may not generalize well across datasets) and does not necessarily involve checking for accuracy of the explanations. On the other hand, faithfulness focuses on the accuracy of individual explanations without such constraints.
2.2 Analyzed Downstream Tasks
Of these, a slim majority of studies on language model consistency investigate well-established NLP tasks. The tasks most commonly analyzed are question answering (QA)Mündler et al. (2024); Raj et al. (2023; 2022); Berglund et al. (2024); Li et al. (2023); Wang et al. (2020); Asai & Hajishirzi (2020a); Liu et al. (2024), summarization West et al. (2024); Cui et al. (2024); Tam et al. (2023); Wang et al. (2020), natural language inference (NLI) Jang et al. (2022); Jang & Lukasiewicz (2023); Camburu et al. (2020); West et al. (2024); Dziri et al. (2019) and reasoning Zhang et al. (2024b); Liu et al. (2024); Chen et al. (2024); Wang et al. (2023) Roughly a third of existing studies do not rely on standard NLP tasks. These studies typically use custom tasks, such as generating continuations of sentences from Wikipedia Mündler et al. (2024). A small number of studies use an approach tailored to specific use cases, e.g. measuring the accuracy of stock price predictions based on textual information such as earnings calls and news articles Yang et al. (2023).
2.3 Size and Availability of Testing Data
The number of testing samples varies quite substantially across different studies, from a few hundred to tens of thousands of samples. One standard approach to creating a test dataset for consistency evaluation is to use one or several of the existing standard benchmarks and modify them according to a set of rules or templates, so that consistency assessment is possible (Jang et al., 2022; Fierro & Søgaard, 2022). Another approach is to enhance existing benchmarks with either human- or LLM-generated annotations (Liu et al., 2023). When evaluating semantic consistency, a common way is to create paraphrases of an existing dataset using automatic parahrasing methods (Bonagiri et al., 2024) and/or human annotators (Elazar et al., 2021). The majority of resulting testing datasets are shared publicly for replicability of experiments, although in some cases the authors rely on the description of the dataset creation process and do not provide access to the actual dataset.
2.4 Evaluated Models
More than two-thirds of the studies on language model consistency in our literature review use generative, transformer-based LMs with decoder-only or encoder-decoder architectures, such as the GPT and OPT series models, BART, and T5 Jang et al. (2022); Li et al. (2023); Jang & Lukasiewicz (2023); West et al. (2024); Berglund et al. (2024); Raj et al. (2023; 2022); Manakul et al. (2023); Mündler et al. (2024); Zhang et al. (2024b); Liu et al. (2024); Zhang et al. (2024a); Cheng et al. (2024); Cui et al. (2024); Tam et al. (2023); Wang et al. (2023); Chen et al. (2024); Wang et al. (2020); Nie et al. (2021); Qin et al. (2021). Model sizes for these studies range from a few billion parameters to hundreds of billions of parameters or more. Many of these studies also consider other types of models. Slightly more than half of the papers we reviewed test proprietary models such as GPT-4 (Li et al., 2023; Jang & Lukasiewicz, 2023; West et al., 2024; Berglund et al., 2024; Mündler et al., 2024; Zhang et al., 2024b; Liu et al., 2024; Zhang et al., 2024a; Cui et al., 2024; Chen et al., 2024; Cheng et al., 2024), whose exact sizes have not been publicly disclosed, but in some cases are rumored to exceed a trillion parameters. About a quarter of the studies Jang et al. (2022); Elazar et al. (2021); Asai & Hajishirzi (2020a); Yang et al. (2023); Nie et al. (2021); Qin et al. (2021) focus on encoder-only, BERT-style models such as BERT, RoBERTa, and ALBERT.
2.5 Evaluation Metrics
Metrics to measure different notions of consistency typically depend on pairwise similarity metrics. They compute intrinsic base metrics such as BERTScore, ROUGE, Entailment, or Contradiction for pairs of outputs given similar inputs and/or context, and aggregate over multiple such pairs. In earlier studies, the base metrics were based on token-matching similarities (Elazar et al., 2021). Later papers graduated to notions of semantic similarity that are robust to syntactic variations that can change the wording or structure of a phrase of text while keeping the meaning the same or similar (Raj et al., 2022; 2023; Rabinovich et al., 2023; Manakul et al., 2023). Aggregation of a metric across a set of pairs is typically done using simple averaging, with the exception of Mündler et al. (2024), which uses sequential aggregation of contradiction scores to detect factual consistency, and (Raj et al., 2023; Kuhn et al., 2023) who use semantic entropy across the full set of outputs.
2.6 Other Important Aspects
In spite of recent attention, certain important aspects of consistency remain underresearched. We specifically call out two such directions. First, recent consistency research tends to focus excessively on consistency in generations at decoding time. In this process it ignores encoder-only models (e.g. those based on BERT and RoBERTa), and how (in)consistent inputs shape the performance on downstream standard NLP tasks such as sentiment prediction. Another direction currently underexplored is adversarial attacks to degrade consistency. While there is extensive research on adversarial robustness (e.g. the AdvGLUE benchmark(Wang et al., 2022)) and jailbreaks, there are very few studies exploring how inconspicuous or subtle manipulation of prompts can lead to inconsistent LLM responses (Lin et al., 2024a), and we do not yet fully understand how much malicious perturbations coupled with slightly different input text can degrade the quality of the downstream output.
The availability of model weights and training datasets—which allow for stronger transparency and reproducibility—aid in investigating the root causes of inconsistency in LLMs. As shown by Lin et al. (2024b), analyzing the internal state of the model can improve the transparency of the model and lay the foundation for mitigating model hallucinations and inconsistencies. Not only closed-weight models, but also unpublished source code and unpublished datasets make it nearly impossible to reproduce some existing publications, making it very hard to verify the claims and findings stated in the publications (Semmelrock et al., 2023).
3 Discussion and Recommendations for Future Research
3.1 Standardization of Consistency Research
In section 2, we reviewed existing consistency research and showed that some types of consistency, such as e.g. transitive or symmetric, and non-English languages are underrepresented in current literature, while other aspects, such as semantic consistency of English-only text-to-text models, are relatively well researched. We also showed that there is very little consensus on the general definition of consistency and the use of evaluation metrics. Due to the lack of standard evaluation processes and benchmarks, the majority of existing work has been performed on individually created datasets that are not always openly shared. This lack of standard evaluation methods makes it difficult to compare the results of individual consistency studies, replicate previous works, and as a result, understand the progress of model development when it comes to consistency evaluation.
3.2 Multilingual Consistency
Similarly to other aspects of NLP research, the overwhelming majority of studies on foundation model consistency are currently performed on English-language data, significantly limiting our understanding of model consistency. To broaden this understanding, more research is needed on both monolingual consistency in non-English languages and on cross-language consistency behaviors.
3.2.1 Consistency in Non-english Languages
There exists a sizable gap between the available amount of training data for English and that available for all other languages Üstün et al. (2024). While more than 7,000 languages are spoken around the world today, an astounding 73% of popular datasets used to train LLMs are primarily or entirely English Longpre et al. (2023). This severe sampling bias in the construction of datasets results in disparities in model performance across languages even on well-studied tasks Lai et al. (2023).
Moreover, inherent differences between languages may significantly influence the consistency of language models trained on them. Structural features such as word order or inflectional morphology can vary in their stability across languages Dediu & Cysouw (2013). These differences can make it more difficult to train models that produce consistent output for certain languages, even when all languages are equally represented in the training data. More research is necessary to understand the effect of linguistic differences and limitations of multilingual training data on consistency in non-English languages.
3.2.2 Cross-lingual Consistency
Cross-lingual consistency refers to whether a model produces compatible or equivalent output when the same query is presented in different languages. Recent work has demonstrated significant challenges in this area. Shen et al. (2024) found that LLMs exhibit inconsistent safety behaviors across languages, with safety guardrails being more easily circumvented in non-English languages. Similarly, Xing et al. (2024) observed that LLMs produce inconsistent factual information when queried about the same knowledge in different languages, suggesting knowledge representation gaps across languages. Qi et al. (2023) examined factual consistency across languages and identified disparities, with languages more dissimilar to English less likely to reflect synthetically inserted factual associations through model editing. Jin et al. (2023) evaluated cross-lingual inconsistency specifically in healthcare questions, finding discrepancies in medical advice across languages. Zhou & Zhang (2024) explored how political biases manifest inconsistently in bilingual models, revealing that models may express different political positions depending on the input language.
These findings collectively highlight a critical gap in current LLM capabilities: the ability to maintain consistent factual information, safety guardrails, and reasoning in different languages. Cross-lingual consistency represents an important direction for future research, especially as LLMs are deployed globally across linguistic boundaries.
3.3 Automatic and Human Evaluation
Most previous studies have used automatic metrics alone to assess the level of consistency in language models. Although automatic evaluation is necessary to ensure objectivity and fast assessment of large data sizes, human evaluation is important to establish an acceptable baseline. It becomes especially important when model consistency is assessed in the context of highly sensitive or subjective culture-based applications (e.g. social appropriateness) or when the automatically measured consistency level is sufficiently high.
The evaluation of consistency poses several unique challenges. Automatic metrics often struggle to capture the full spectrum of consistency types (factual, logical, semantic), while human evaluation suffers from subjectivity and cognitive biases. The contextual nature of consistency requires evaluation across multiple responses, different phrasings, and various contexts, making comprehensive assessment computationally expensive and logistically challenging. Further complicating matters, consistency evaluation interacts with other dimensions such as factuality, helpfulness, and safety—a model may be internally consistent but factually incorrect, or it may sacrifice consistency to maintain safety.
Interestingly, there is no publicly available benchmark yet to measure all different aspects of consistency in language models across diverse tasks. Future work should focus on developing high-quality benchmarks that address the above challenges.
3.4 The Harms and Benefits of Consistency
Inconsistent outputs from language models can lead users to receiving conflicting or incorrect information, which is particularly problematic in scenarios where factual accuracy is crucial (Tam et al., 2023; Wang et al., 2024)—such as in medical, legal, or financial contexts—especially when such information is used for decision-making. In critical systems, such as autonomous vehicles or medical diagnosis support, inconsistent responses can lead to critical safety risks. In less critical applications, inconsistent responses lead to poor user experience, causing frustration and reducing overall utility of the system (Lazar et al., 2023; van Bergen et al., 2024; Zhang et al., 2024a).
In addition to the above problems, inconsistent responses can reflect and magnify the underlying societal biases and stereotypes in training data, leading to potentially discriminatory outcomes for certain user groups, amplifying unfair use of the models, and causing representational harm, including demeaning social groups (Blodgett et al., 2020).
While inconsistency causes certain risks and potential harms to the users of language models in real-world applications, it is also important to understand its potential value and advantages in specific settings and situations. First, some degree of inconsistency can lead to more diverse and creative output, which can be valuable in tasks requiring originality or brainstorming. As such, inconsistency might reflect the ability of a model to adapt to different contexts or user needs, potentially providing more personalized responses.
Second, inconsistent outputs of a model prompt users to engage more critically with the content and seek additional verification. This can be beneficial in educational applications, as long as the level of possible model inconsistency is known and carefully controlled.
Finally, inconsistencies can help identify areas where the model needs improvement, driving further research and development in AI.
3.5 Methods to Improve Consistency
Despite the importance, surprisingly few approaches have emerged that actually increase the consistency of language models. Current proposals to do so fall into two narrow categories. The first uses fine-tuning techniques to improve consistency between multiple generations from a language model when supplied with same or similar outputs. To this end, (Elazar et al., 2021) used a custom loss function, Raj et al. (2025) used knowledge distillation from more consistent teacher models, and Raj et al. (2025); Zhao et al. (2024) utilized synthetic datasets of groups of consistent input-outputs. The second approach attempts to improve self-consistency, i.e. consistency between the reasoning process a model goes through and the final answer it produces Deng et al. (2023); Wang et al. (2023); Wei et al. (2022).
While these methods show promise, they primarily address symptoms rather than the fundamental causes of inconsistency. There remains a critical need for research investigating the structural basis of consistency in language models’ representational spaces, consistency-oriented pretraining objectives, and architectures designed to maintain consistency across diverse contexts. Such foundational approaches could potentially eliminate the trade-offs between consistency and other valuable properties like creativity or adaptability that the above methods tend to exhibit.
4 Conclusion and Call to Action
In this paper, we provide a review of consistency in language models, highlighting challenges, current approaches, and future directions in this critical area of research. We have shown that despite significant advances, the field lacks standardized definitions, evaluation metrics, benchmarks, and mitigation techniques. This lack of standardization hampers progress and makes it difficult to compare the results between studies.
We specifically call on the research community to address several key challenges: (1) developing standardized definitions and taxonomies of consistency types; (2) creating comprehensive, multilingual, and cross-lingual benchmarks for consistency evaluation; (3) establishing robust evaluation protocols that combine automatic metrics with human evaluation; (4) investigating the relationship between consistency and other important properties such as factuality, safety, and helpfulness; and (5) developing efficient methods to enhance consistency without sacrificing other beneficial model capabilities.
To this end, we emphasize the need for interdisciplinary collaboration, bringing together perspectives from linguistics, psychology, philosophy, and ethics to better understand the multifaceted nature of consistency in human and machine language use. By addressing these challenges collectively, we can move toward language models that exhibit more reliable, trustworthy, and human-aligned behavior across diverse contexts and applications.
Appendix A Appendix: Multimodal Consistency
Until 2022, every consistency study was analyzing robustness of language models to various text perturbations or to semantically equivalent texts only. Starting in 2022, some interest in non-textual modalities started to appear that comes primarily from text-to-image model analysis. For example, Tan et al. (2022) explores the challenge of generating consistent and high-quality images from given texts in the task of visual-language understanding and highlights the need to design a better text-image consistency metric, a problem that remains under-explored in the community. In their study, Tan et al. (2022) present a novel CLIP-based metric named Semantic Similarity Distance (SSD) that leads to significantly better text-image consistency while maintaining decent image quality. The attempts to quantify the consistency in text-to-image models are continued by Berglund et al. (2024), which proposes a novel semantic consistency score for image generation that has strong agreement with human annotators. Recently, there was an attempt to evaluate the understanding capability of generative models in both language and vision domains West et al. (2024). West et al. (2024) conducted interrogative evaluation of image understanding models via visual question answering in an open-ended setting. They investigated whether the models produce consistent output when interrogated about the content of their generated image, and figured out that although models can outperform humans in generation, they regularly show evidence of inconsistency between their generation and understanding performance.